A survey asks a random sample of 1500 adults in Ohio if they support an increase in the state sales tax from 5% to 6%, with the additional revenue going to education. Let P denote the proportion in the sample who say they support the increase. Suppose that 54% of all adults in Ohio support the increase. The standard deviation of the sampling distribution is...? Round your answer to four decimal places.
Answer #1

Solutioin

\mathbf{given:}\;p=0.54,n=1500

standard\; deviation \;of \;the \;sampling\; distribution\; is=\sqrt{\frac{p*(1-p)}{n}}

\\standard\; deviation \;of \;the \;sampling\; distribution\; is=\sqrt{\frac{0.54*(1-0.54)}{1500}}

\\standard\; deviation \;of \;the \;sampling\; distribution\; is=\mathbf{{\color{Red} 0.0129}}